Suppose the short run production function can be represented by Q=60000L0^2_1000b^3. Then determine



The level of labor employment that maximizes the level of out put

1
Expert's answer
2022-01-12T08:36:20-0500

The level of labor employment that maximizes the level of output is:

"Q'(L) = 120,000L - 3000L^2 = 0,"

3,000L(40 - L) = 0,

L = 0 (is not acceptable), L = 40 units.